# Task-Agnostic Continual Prompt Tuning with Gradient-Based Selection and Decoding

### v1

@@ 6 Conclusion
We proposed GRID, a unified framework for scalable and forgetting-resilient prompt-based continual learning. GRID addresses two critical challenges in continual prompt tuning: latent forgetting under task-agnostic inference and unbounded prompt growth as task sequences increase. By combining task-aware decoding with a novel gradient-based prompt selection and compression strategy, GRID enables efficient prompt reuse and effective task retention. Our method significantly improves backward knowledge retention, reducing the number of forgotten tasks by up to 91.8% in certain settings, while maintaining competitive forward transfer. Extensive experiments across short-sequence, long-sequence, and negative transfer benchmarks, using both T5 and Flan-T5 backbones, demonstrate that GRID achieves state-of-the-art performance in both accuracy and forgetting metrics, making it a practical solution for realistic, long-horizon continual learning.
@@ Appendix A Algorithmic Details
Appendix B Further Implementation Details and Additonal Results
@@ B.1 Datasets
Table 8: Overview of the 15 datasets used in our CL experiments, including their evaluation metrics. Datasets from CL benchmark [52], GLUE [39], and SuperGLUE [40] were utilized, along with the IMDB movie reviews dataset. For tasks with two evaluation metrics, we report the average as the final performance measure.
@@ 15. IMDB Other sentiment analysis movie reviews accuracy
Short-Sequence CL Setting:
We evaluate our method on four widely-used text-classification datasets introduced by [52]: Amazon Reviews, DBpedia, Yahoo Answers, and AG News. We follow the few-shot continual learning setting outlined by [25] and [27], evaluating across three distinct task sequences (Orders 1–3) [27].
Long-Sequence CL Setting:
Following the approach of [27], we consider a total of 15 distinct tasks. These tasks include the four previously described in the short-sequence CL benchmark, four tasks from the GLUE benchmark (MNLI, QQP, RTE, SST2) [39], five tasks from the SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ) [40], as well as YELP reviews [52] and the IMDB movie reviews dataset [21]. We evaluate our methods across six different task sequence orders. Three of these sequences follow standard long-sequence continual learning orders that have been used in prior works (e.g., Orders 4–6)[27, 13]. We additionally propose three novel task orders: (1) a Order $A$: easy-to-hard order, where tasks are categorically grouped and simpler tasks such as binary sentiment classification are introduced first, followed by more complex reasoning tasks like MultiRC and COPA; (2) a Order $B$: hard-to-easy order that reverses this progression; and (3) a Order $C$: mixed order that randomly interleaves tasks from different categories and difficulties. These variations allow us to assess how the order of task presentation and semantic similarity affect forgetting, transfer, and model generalization in long-horizon continual learning scenarios.
B.2 Limitations and Future Work
While GRID demonstrates strong improvements in backward transfer and scalability, it has two primary limitations. First, our evaluation is limited to sequences of 15 tasks. To fully assess the scalability of our method, future work should test it on significantly longer task streams. Second, the current GRID framework—specifically the task-aware decoding component—is designed for encoder-decoder architectures. Although our gradient-based prompt selection strategy is model-agnostic and can be applied broadly, the full framework has not yet been extended to encoder-only or decoder-only models. Exploring how to adapt GRID’s decoding mechanism to these architectures is a promising direction for future research.
Table 9: Task sequence orders used in our CL experiments. Orders 1-3 represent standard short-sequence CL benchmarks, while Orders 4-6 correspond to long-sequence CL benchmarks. Orders A-C are our proposed custom sequences, with A representing an easy-to-hard order, B representing a hard-to-easy order, and C representing a mixed order. NT1, NT2, and NT3 represent different task evaluation orders designed to examine negative transfer across various task sequences.
@@ B.2 Limitations and Future Work
1 db $\rightarrow$ amazon $\rightarrow$ yahoo $\rightarrow$ ag
2 db $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo
3 yahoo $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ db
4 mnli $\rightarrow$ cb $\rightarrow$ wic $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ boolq $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ multirc $\rightarrow$ yahoo
5 multirc $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ yahoo
6 yelp $\rightarrow$ amazon $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ multirc $\rightarrow$ boolq $\rightarrow$ wic
A sst2 $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ dbpedia $\rightarrow$ mnli $\rightarrow$ rte $\rightarrow$ cb $\rightarrow$ qqp $\rightarrow$ copa $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ multirc
B multirc $\rightarrow$ wic $\rightarrow$ boolq $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ cb $\rightarrow$ rte $\rightarrow$ mnli $\rightarrow$ dbpedia $\rightarrow$ yahoo $\rightarrow$ ag $\rightarrow$ amazon $\rightarrow$ yelp $\rightarrow$ imdb $\rightarrow$ sst2
C sst2 $\rightarrow$ copa $\rightarrow$ ag $\rightarrow$ imdb $\rightarrow$ mnli $\rightarrow$ yahoo $\rightarrow$ rte $\rightarrow$ yelp $\rightarrow$ qqp $\rightarrow$ cb $\rightarrow$ amazon $\rightarrow$ dbpedia $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ multirc
@@
Table 10: Accuracy for different datasets with varying sample sizes.
Dataset No. of samples/k Accuracy
dbpedia 20 0.5203
200 0.9674
1k 0.9880
1.5k 0.9892
amazon 20 0.0000
200 0.0000
1k 0.5136
1.5k 0.5387
ag 20 0.0000
200 0.8180
1k 0.8900
1.5k 0.8933
@@
Table 11: Impact of prompt length on T5-large performance across Orders 4–6. We observe a trade-off: longer prompts improve BWT (less forgetting) but increase the number of forgotten tasks and can degrade accuracy. Optimal performance is generally observed at a moderate prompt length (e.g., 10).
Prompt Length Order 4 Order 5 Order 6
@@
Table 12: Average test accuracy across long-sequence task orderings (Order 4–6, A–C) for T5 and Flan-T5 models. Our method consistently outperforms Progressive Prompts in nearly all configurations.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg
@@
Table 15: Backward Transfer (BWT) scores on general task orderings (Order 1–3) comparing our method with Progressive Prompts across multiple models. Our method consistently achieves less negative BWT, indicating improved retention of prior tasks across varying task sequences.
Model Setting Order 1 Order 2 Order 3 Avg
T5-small Baseline -0.5284 -0.5214 -0.5899 -0.5466
Ours -0.3102 -0.3010 -0.2471 -0.2861
FT5-small Baseline -0.6116 -0.5871 -0.6462 -0.6150
Ours -0.3590 -0.3254 -0.3265 -0.3370
T5-base Baseline -0.6499 -0.6080 -0.6406 -0.6328
Ours -0.4039 -0.3494 -0.2755 -0.3430
FT5-base Baseline -0.6594 -0.6457 -0.6991 -0.6681
Ours -0.4187 -0.3722 -0.2679 -0.3529
T5-large Baseline -0.6407 -0.6311 -0.7081 -0.6600
Ours -0.4221 -0.3705 -0.3193 -0.3706
FT5-large Baseline -0.6808 -0.6797 -0.7510 -0.7038
Ours -0.4863 -0.4099 -0.3345 -0.4103
Table 16: Average test accuracy on short-sequence continual learning experiments (Order 1–3). Our method performs competitively across all model variants, matching or outperforming the baseline, especially for larger models like FT5-base, while not applying prompt compression due to the limited number of tasks.
Model Setting Order 1 Order 2 Order 3 Avg
T5-small Baseline 65.29 66.11 67.55 66.32
Ours 64.43 64.94 66.05 65.14
FT5-small Baseline 72.15 70.04 72.75 71.65
Ours 72.79 70.90 72.81 72.17
T5-base Baseline 75.70 73.30 72.55 73.85
Ours 75.25 75.65 74.56 75.15
FT5-base Baseline 76.87 76.92 77.12 76.97
Ours 77.59 77.56 76.09 77.08
T5-large Baseline 75.13 77.32 77.73 76.73
Ours 74.84 74.41 74.55 74.60
FT5-large Baseline 80.40 81.23 81.20 80.94
Ours 80.29 80.95 80.61 80.62
Table 17: Forward Transfer (FWT) scores across various task orderings (Order 4–6, A–C) for T5 and Flan-T5 models. Higher scores indicate better transfer to future tasks. Our method achieves comparable or improved FWT in most cases, demonstrating that reducing forgetting does not come at the cost of forward transfer.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg
@@
Table 19: Forward Transfer (FWT) scores on short-sequence tasks (Orders 1–3) comparing our method with Progressive Prompts for T5-small and FT5-small. While our method prioritizes backward retention, it maintains competitive FWT performance, confirming that prompt compression does not significantly hinder future task generalization.
Model Setting Order 1 Order 2 Order 3 Avg
T5-small Baseline -5.93 -5.39 -3.73 -5.02
Ours -6.81 -6.95 -6.39 -6.72
FT5-small Baseline -2.12 -5.07 -1.69 -2.96
Ours -2.49 -4.95 -2.66 -3.37
T5-base Baseline 0.72 -2.45 -3.56 -1.76
Ours 0.93 1.57 -0.02 0.83
FT5-base Baseline -2.46 -2.45 -2.23 -2.38
Ours -1.94 -2.07 -3.58 -2.53
T5-large Baseline -3.41 -0.69 -0.02 -1.37
Ours -4.05 -4.53 -4.03 -4.20
FT5-large Baseline -1.01 0.20 -0.16 -0.32
Ours -0.26 0.71 -0.23 0.07
(A) Baseline
(B) Ours
(c) B-A
Figure 4: Heatmaps of backward transfer scores on previous tasks for Order 5. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 5: Per-task BWT comparison between our method (blue) and the baseline (red) for Order 5. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 6: Heatmaps of backward transfer scores on previous tasks for Order 6. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 7: Per-task BWT comparison between our method (blue) and the baseline (red) for Order 6. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 8: Heatmaps of backward transfer scores on previous tasks for Order A. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 9: Per-task BWT comparison between our method (blue) and the baseline (red) for Order A. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 10: Heatmaps of backward transfer scores on previous tasks for Order B. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 11: Per-task BWT comparison between our method (blue) and the baseline (red) for Order B. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 12: Heatmaps of backward transfer scores on previous tasks for Order C. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 13: Per-task BWT comparison between our method (blue) and the baseline (red) for Order C. Positive bars indicate improved retention of prior tasks.
Table 20: Summary of the results with the T5-large model on the short-sequence benchmark. Average accuracy after training on the last task is reported. All results are averaged over 3 runs. For the T5 model, we follow few-shot CL settings as in [25]
Method Order 1 Order 2 Order 3 avg
Per-task $70.0$ $70.0$ $70.0$ $70.0$
Continual Finetune $18.9$ $24.9$ $41.7$ $28.5$
Data Replay $35.4$ $37.1$ $41.5$ $38.0$
EWC $39.0$ $38.0$ $44.8$ $40.6$
LFPT5 $47.6$ $52.6$ $57.9$ $52.7$
CODA-Prompt $77.0$ $80.2$ $75.9$ $77.7$
ProgPrompt $74.1$ $74.2$ $75.3$ $74.5$
SHLPT $80.2$ $79.7$ $80.9$ $80.1$
Ours* 74.84 74.41 74.55 74.60
B.3 Implementation Details
In our experiments, we focus on encoder-decoder transformer-based architecture, adopting both the original T5[26] and Flan-T5[7] as our backbone architectures. To study the impact of model size, we perform evaluations using the small, base, and large variants. This setup allows us to assess the scalability of our method across a range of model capacities.
Experimental Setting
In all our experiments with T5 and Flan-T5 backbones, we fix the prompt length to 10 tokens, train for 10 epochs, and use 1k representative samples per class. [^3] We could not experiment with 300 epochs and 20 samples per class (as in ProgPrompt [27]) due to resource constraints. This configuration is applied uniformly across the small, base, and large variants of both model families, ensuring a fair yet efficient comparison to prior work such as ProgPrompt [27]. All experiments were performed on a single A100 40GB GPU with batch size set to 8 for all experiments.

### v3

@@ 6 Conclusion
In this work, we introduce GRID, a unified framework that renders prompt-based continual learning both scalable and resilient to forgetting. Unlike prior approaches that depend on task-aware inference or accumulate ever-growing prompt pools, GRID tackles two key challenges: latent forgetting under task-agnostic inference and the inefficiency of unbounded prompt memory. It integrates constrained decoding with gradient-guided prompt selection and compression, enabling consistent label generation, compact memory usage, and enhanced knowledge retention. Our method substantially improves backward transfer, reducing forgotten tasks by a large margin, while maintaining competitive forward transfer.
Future work could extend GRID to also enable positive backward knowledge transfer, allowing new tasks to refine earlier prompts and further boost their performance. Additional directions include scaling to much longer task streams, applying the framework to other types of foundation models.
Supplementary Materials
@@ Appendix A Algorithmic Details
Appendix B Further Implementation Details
@@ B.1 Datasets
Table 9: Overview of the 15 datasets used in our CL experiments, including their evaluation metrics. Datasets from CL benchmark (Zhang et al., 2015), GLUE (Wang et al., 2018), and SuperGLUE (Wang et al., 2019) were utilized, along with the IMDB movie reviews dataset. For tasks with two evaluation metrics, we report the average as the final performance measure.
@@ 15. IMDB Other sentiment analysis movie reviews accuracy
Table 10: Task sequence orders used in our CL experiments. Order L1-L3 correspond to long-sequence CL benchmarks. Order L4-L6 are our proposed custom sequences, with L4 representing an easy-to-hard order, L5 representing a hard-to-easy order, and L6 representing a mixed order. NT1, NT2, and NT3 represent different task evaluation orders designed to examine negative transfer across various task sequences.
@@
L1 mnli $\rightarrow$ cb $\rightarrow$ wic $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ boolq $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ multirc $\rightarrow$ yahoo
L2 multirc $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ yahoo
L3 yelp $\rightarrow$ amazon $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ multirc $\rightarrow$ boolq $\rightarrow$ wic
L4 sst2 $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ dbpedia $\rightarrow$ mnli $\rightarrow$ rte $\rightarrow$ cb $\rightarrow$ qqp $\rightarrow$ copa $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ multirc
L5 multirc $\rightarrow$ wic $\rightarrow$ boolq $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ cb $\rightarrow$ rte $\rightarrow$ mnli $\rightarrow$ dbpedia $\rightarrow$ yahoo $\rightarrow$ ag $\rightarrow$ amazon $\rightarrow$ yelp $\rightarrow$ imdb $\rightarrow$ sst2
L6 sst2 $\rightarrow$ copa $\rightarrow$ ag $\rightarrow$ imdb $\rightarrow$ mnli $\rightarrow$ yahoo $\rightarrow$ rte $\rightarrow$ yelp $\rightarrow$ qqp $\rightarrow$ cb $\rightarrow$ amazon $\rightarrow$ dbpedia $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ multirc
@@
B.2 Task Sequence Orders
Long-Sequence CL Setting:
Following the approach of (Razdaibiedina et al., 2023), we consider a total of 15 distinct tasks. These consist of five datasets from the CL benchmark (Zhang et al., 2015): AG News (topic classification), Amazon Reviews (sentiment analysis), Yelp Reviews (sentiment analysis), DBpedia (Wikipedia text classification), Yahoo Answers (Q&A classification), four tasks from the GLUE benchmark (MNLI, QQP, RTE, SST2) (Wang et al., 2018), five tasks from the SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ) (Wang et al., 2019), and the IMDB movie reviews dataset for sentiment analysis (Maas et al., 2011). We evaluate our methods across six different task sequence orders. Three of these sequences follow standard long-sequence continual learning orders that have been used in prior works (e.g., L1–L3) (Razdaibiedina et al., 2023; Guo et al., 2024). We additionally propose three novel task orders: (1) Order L4: easy-to-hard, where tasks are grouped by category and simpler tasks such as binary sentiment classification are introduced first, followed by more complex reasoning tasks like MultiRC and COPA; (2) Order L5: hard-to-easy, which reverses this progression; and (3) Order L6: mixed, where tasks from different categories and difficulty levels are randomly interleaved. These variations allow us to assess how task presentation order and semantic similarity affect forgetting, transfer, and model generalization in long-horizon continual learning scenarios.
Negative Transfer Benchmark. To evaluate the robustness of our approach against negative transfer (Pan, 2010), we introduce a Negative Transfer Benchmark (NT1, NT2, NT3) inspired by the SHLPT framework (Wu et al., 2024). Specifically, we identify task pairs that exhibit negative transfer based on the analysis provided in the SHLPT. Using these insights, we create three task sequences that are likely to cause negative transfer, meaning earlier tasks make it harder to learn later ones. This helps us test how well our method handles such challenges.
B.3 Implementation Details
In our experiments, we focus on encoder-decoder transformer-based architecture, adopting both the original T5(Raffel et al., 2020) and Flan-T5(Chung et al., 2022) as our backbone architectures. To study the impact of model size, we perform evaluations using the small, base, large, and 3B variants. This setup allows us to assess the scalability of our method across a range of model capacities.
B.4 Experiment Details
In all our experiments with T5 and Flan-T5 backbones, we fix the prompt length to 10 tokens, train for 10 epochs, and use 1k representative samples per class. [^3] We could not experiment with 300 epochs and 20 samples per class (as in ProgPrompt (Razdaibiedina et al., 2023)) due to resource constraints. This configuration is applied uniformly across the small, base, and large variants of both model families, ensuring a fair yet efficient comparison to prior work such as ProgPrompt (Razdaibiedina et al., 2023).
Appendix C Additional Results
C.1 Order and BWT Analysis
One critical observation from Table 11 and Table 14 is that Order L5 consistently outperforms Order L4 in both BWT scores and the number of forgotten tasks. Specifically, for the same set of tasks, Order L5 shows less negative BWT scores, indicating reduced forgetting. Additionally, the number of forgotten tasks is lower in Order L5, further suggesting that harder tasks at the beginning of training help the model retain earlier knowledge more effectively. This observation highlights the benefit of starting with more difficult tasks, as it leads to better long-term retention and improved overall performance in continual learning. One also notes that Order L6 performs better than Order L4 but slightly less effectively than Order L5, indicating that a mixed task sequence provides a balanced approach to task progression, offering a compromise between the benefits of easy-to-hard and hard-to-easy learning.
Forgotten Task Count. We further analyze the number of forgotten tasks across various settings in Table 14. Our method significantly reduces the number of forgotten tasks compared to the Progressive Prompts baseline. For instance, in the case of T5-small, the number of forgotten tasks drops from as high as 93 in some orders to as low as 7 under our method. Similar improvements are observed for T5-base, where the count decreases from 90 to just 6 in the best case, and for T5-large, where it reduces from 87 to only 5. These improvements are consistent across all task orderings, including structured sequences such as Order L4 (easy-to-hard), Order L5 (hard-to-easy), and Order L6 (mixed).
Table 11: Backward Transfer (BWT) comparison between our method and ProgPrompt across multiple task orderings (Order L1-L6) for T5 and Flan-T5 models. Less negative scores indicate reduced forgetting. Our method consistently improves BWT across all model variants.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg Imp. (%)
Ours -0.1402 -0.1894 -0.2504 -0.3141 -0.2734 -0.3006 -0.2447
Ours -0.0518 -0.1449 -0.2106 -0.2129 -0.1517 -0.2396 -0.1686
Ours -0.3060 -0.2404 -0.3549 -0.2893 -0.3439 -0.3259 -0.3101
Ours -0.2616 -0.3243 -0.3319 -0.3245 -0.3240 -0.2743 -0.3067
Ours -0.3243 -0.3336 -0.3979 -0.3912 -0.2956 -0.3512 -0.3490
Ours -0.3115 -0.3656 -0.3544 -0.3514 -0.3282 -0.3540 -0.3442
Appendix D Limitations and Future Work
While GRID demonstrates strong improvements in backward transfer and scalability, it has two primary limitations. First, our evaluation is limited to sequences of 15 tasks. To fully assess the scalability of our method, future work should test it on significantly longer task streams. Second, the current GRID framework, specifically the constrained decoding component is designed for encoder-decoder architectures. Although our gradient-based prompt selection strategy is model-agnostic and can be applied broadly, the full framework has not yet been extended to other foundational models. Exploring how to adapt GRID’s decoding mechanism to these architectures is a promising direction for future research.
@@ Appendix D Limitations and Future Work
Table 12: Impact of prompt length on T5-large performance across Order L1-L3. We observe a trade-off: longer prompts improve BWT (less forgetting) but increase the number of forgotten tasks and can degrade accuracy. Optimal performance is generally observed at a moderate prompt length (e.g., 10).
Prompt Length L1 L2 L3
@@
Table 13: Average test accuracy across long-sequence task orderings (Order L1-L6) for T5 and Flan-T5 models. Our method consistently outperforms Progressive Prompts in nearly all configurations.
Model Setting L1 L2 L3 L4 L5 L6 Avg
@@
Table 14: Comparison of forgotten task counts between Progressive Prompts (Baseline) and our method for different models across various task orders (order L1–L6).
Model Setting L1 L2 L3 L4 L5 L6 Average Improvement (%)
T5-small ProgPrompt 85 85 88 90 84 93 87.5 76.8%
Ours 14 7 28 36 16 21 20.3
FT5-small ProgPrompt 93 95 95 94 97 93 94.5 76.7%
Ours 17 14 28 26 20 27 22.0
T5-base ProgPrompt 88 82 76 90 83 88 84.5 80.9%
Ours 15 6 19 29 13 15 16.2
FT5-base ProgPrompt 94 93 93 93 93 93 93.2 83.0%
Ours 10 10 15 25 14 21 15.8
T5-large ProgPrompt 77 72 80 87 71 85 78.7 82.4%
Ours 11 5 18 26 10 13 13.8
FT5-large ProgPrompt 92 93 93 93 91 93 92.5 79.8%
Ours 15 10 24 24 14 25 18.7
@@
Table 17: Forward Transfer (FWT) scores across various task orderings (Order L1-L6) for T5 and Flan-T5 models. Higher scores indicate better transfer to future tasks. Our method achieves comparable or improved FWT in most cases, demonstrating that reducing forgetting does not come at the cost of forward transfer.
Model Setting L1 L2 L3 L4 L5 L6 Avg
@@
Figure 3: Per-task BWT comparison between our method (blue) and the baseline (red) for Order L1. Positive bars indicate improved retention of prior tasks. Our method shows significant BWT gains on several tasks (e.g., copa, wic, yahoo), demonstrating its effectiveness in mitigating forgetting across diverse task types.
(A) Baseline
(B) Ours
(c) B-A
Figure 4: Heatmaps of backward transfer scores on previous tasks for Order L2. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 5: Per-task BWT comparison between our method (blue) and the baseline (red) for Order L2. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 6: Heatmaps of backward transfer scores on previous tasks for Order L3. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 7: Per-task BWT comparison between our method (blue) and the baseline (red) for Order L3. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 8: Heatmaps of backward transfer scores on previous tasks for Order L4. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 9: Per-task BWT comparison between our method (blue) and the baseline (red) for Order L4. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 10: Heatmaps of backward transfer scores on previous tasks for Order L5. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 11: Per-task BWT comparison between our method (blue) and the baseline (red) for Order L5. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 12: Heatmaps of backward transfer scores on previous tasks for Order L6. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 13: Per-task BWT comparison between our method (blue) and the baseline (red) for Order L6. Positive bars indicate improved retention of prior tasks.